# GenIAS: Generator for Instantiating Anomalies in time Series

### v1

@@ 3.2.5. Enhanced KL Divergence Loss
The standard KL divergence loss is typically used to regularize the latent space by aligning the approximate posterior distribution $q(z|\mathbf{X})$, parameterized by $\mu$ and $\sigma^{2}$, with a standard Gaussian prior $p(z)$, defined as $\mathcal{N}(0,1)$. While effective for many tasks, this standard approach may fail to create the tight and compact latent representations necessary for generating anomalies that are distinguishable from normal samples.
To address this, we propose an enhanced KL divergence loss by incorporating a tunable prior variance, $\sigma_{\text{prior}}^{2}$, where $\sigma_{\text{prior}}<1$. This modification enforces tighter representations of normal samples in the latent space. The modified KL divergence loss is defined as:
(5) $\mathcal{L}_{\text{en-KL}}=-\frac{1}{2}\sum_{i=1}^{|\mathbf{\mathcal{D}}|}\sum_{j=1}^{L}\Big{[}1+\log(\sigma^{2}_{ij})-\mu_{ij}^{2}-\frac{\sigma^{2}_{ij}}{\sigma_{\text{prior}}^{2}}+2\log(\sigma_{\text{prior}})\Big{]},$
@@
Where $\mu_{ij}$ and $\sigma^{2}_{ij}$ are the mean and variance of the latent distribution for the $j$-th dimension of the $i$-th sample, and $\sigma_{\text{prior}}$ is the prior’s standard deviation.
Theorem 1 (Enhanced Separation Between Normals and Anomalies in Latent Space).
@@
The proof is provided in Appendix D.3. We also illustrate in Figure 8 in the Appendix how our enhanced KL loss leads to the separation of reconstruction error in the raw space.
The benefits of the enhanced KL divergence loss are as follows:
  • Controlled spread: $\sigma_{\text{prior}}$ determines the spread of normal samples in the latent space. Smaller values enforce tighter clustering, leading to compact representations while allowing limited diversity for variability. (See Lemma 1)
  • Regularization: The KL loss with $\sigma_{\text{prior}}$ ensures a well-regularized latent space, reducing overfitting and promoting smooth sampling.
  • Improved normal sample representation: By constraining the latent space, normal samples occupy a distinct, dense region, enhancing their separability from anomalies. (See Lemma 2 and Theorem 1)
  • Enhanced interpretability: The structured latent space shaped by $\sigma_{\text{prior}}$ provides consistent and interpretable embeddings for both normal and anomalous data.
3.2.6. Total Loss
The total loss integrates all the above components, each weighted by a corresponding hyperparameter:
@@ 3.2.6. Total Loss
(6) $\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{en-KL}},$
@@
Where $\alpha$, $\beta$, $\gamma$, and $\zeta$ are the weights for the reconstruction, perturbation, zero-perturbation, and KL divergence losses, respectively.
3.3. Deviation Patching Approach
After training the GenIAS model in the first step, the anomaly injection process uses the pretrained model to perturb the latent space representation of time series windows ($\mathbf{X}$) and generate realistic anomalous samples ($\tilde{\mathbf{X}}$). This step consists of two phases: (1) anomaly injection via latent space perturbation and (2) patching the injected anomalies using a deviation-based approach, defined by a scaling factor $\tau$, to ensure the generated samples are realistic and aligned with the domain requirements. The injection and patching process is applied independently for each dimension ($d$) of the MTS, and more details are provided in Algorithm 2.
@@ 3.3. Deviation Patching Approach
Figure 3. Effect of deviation-based patching with different thresholds (0.4, 0.2, 0.05) on generated anomalous windows across various entities from the MSL, SMAP, SMD, Yahoo, and KPI datasets.
4. Experiments
@@
We conduct extensive evaluations of GenIAS’s performance using a comprehensive selection of datasets and baselines. Our evaluation consists of two parts: (1) Generation Evaluation, where we compare GenIAS against other anomaly injection/generation models to assess its capability in generating realistic and diverse anomalies, and (2) Case Study – TSAD Evaluation, where we benchmark GenIAS against seventeen competing methods using nine real-world datasets covering both MTS and UTS. Specifically, to study the effectiveness of GenIAS’s anomaly generation in TSAD, we integrate GenIAS with the anomaly detector architecture from CARLA to implement a GenIAS-enhanced TSAD model (i.e., substituting CARLA’s anomaly generator with GenIAS while keeping the rest of the framework identical). We choose CARLA as our base detector due to its SOTA performance, which enables direct comparison with the strongest generation-based TSAD method. Through this comprehensive evaluation, we provide a thorough analysis of GenIAS’s capabilities in both anomaly generation and TSAD tasks.
Table 1. An overview of dataset statistics.
@@
Benchmark # datasets # dims Train size Test size ARMSL (Hundman et al., 2018) 27 55 58,317 73,729 10.72% SMAP (Hundman et al., 2018) 55 25 140,825 444,035 13.13% SMD (Su et al., 2019) 28 38 708,405 708,420 4.16% SWaT (Mathur and Tippenhauer, 2016) 1 51 496,800 449,919 12.33% GECCO (Moritz et al., 2018) 1 9 69,260 69,261 1.1% SWAN (Angryk et al., 2020) 1 38 60,000 60,000 32.6% UCR (Wu and Keogh, 2021) 205 1 2,238,349 6,143,541 0.6% Yahoo-A1 (Laptev et al., 2015) 67 1 46,667 46,717 1.76% KPI (aio, 2018) 29 1 1,048,576 2,918,847 1.87%
4.1. Benchmark Datasets
@@
GenIAS is thoroughly evaluated on nine widely adopted real-world benchmark datasets for TSAD. Their key statistics are reported in Table 1. In particular, our selection includes six MTS datasets (MSL, SMAP, SMD, SWaT, GECCO, and SWAN) and three UTS datasets (UCR, Yahool-A1 and KPI), containing diverse anomaly types and a variety of anomaly ratios, thereby providing a comprehensive evaluation of the performance of our proposed methods and the baselines. Detailed dataset descriptions can be found in Appendix A.
4.2. Baselines
@@
GenIAS is compared with seventeen relevant and competitive baseline methods that leverage a wide array of anomaly detection approaches, which can be categorised into four broad categories: classic anomaly detection methods, positive-unlabeled (P-U) methods, unsupervised deep TSAD methods that do not leverage anomaly generation (US-WI), and unsupervised TSAD methods that employ anomaly generation (US-I).
@@
Specifically, three classic general AD methods - OCSVM (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008) - are included to establish a base performance for demonstrating the advantages of deep learning TSAD approaches. LSTM-VAE (Park et al., 2018), a deep learning TSAD method that utilises labeled anomalies, is included to examine the effects of leveraging labelled anomalies for TSAD. US-WI is another relevant category, which includes a variety of unsupervised TSAD baselines that do not leverage anomaly generation. Donut (Xu et al., 2018), OmniAnomaly (Su et al., 2019), AnomalyTransformer (Xu et al., 2021), and TranAD (Tuli et al., 2022b) use reconstruction error to detect anomalies. THOC (Shen et al., 2020b) and TimesNet (Wu et al., 2023) treat TSAD as a forecasting problem. MTAD-GAT (Zhao et al., 2020) integrates graph attention mechanisms to capture both temporal and multivariate dependencies. Representation learning methods like TS2Vec (Yue et al., 2022) and DCdetector (Yang et al., 2023) focus on establishing robust and anomaly-discriminative representations. US-I is the most relevant category, which allows us to compare GenIAS with existing generation methods for unsupervised TSAD. We include four recent unsupervised TSAD methods that leverage various anomaly generation approaches: NCAD (Carmona et al., 2022), CutAddPaste (Wang et al., 2024), COUTA (Xu et al., 2024), and CARLA (Darban et al., 2025).
4.3. Evaluation Metrics
@@
We adopt seven evaluation metrics to assess GenIAS’s performance in terms of generation quality and TSAD performance.
@@
Generation quality. To quantitatively examine the quality of anomaly generation, we introduce two key metrics: Anomalous Representation Proximity (ARP), which assesses realism, and Entropy-Based Diversity Index (EDI), which characterizes the diversity of the generated anomalies, motivated by the idea in AnomalyDiffusion (Hu et al., 2024). Due to page limits, the definitions and discussions of these metrics are provided in the Appendix E. In general, ARP is the primary metric as it measures the closeness of the generated anomalies to the actual anomalies, see Appendix E.1. EDI is complementary to ARP, as diversity should only be examined under the condition that the generated anomalies can illustrate the actual anomalies, see Appendix E.2.
@@
TSAD evaluation. We consider five widely-used, complementary metrics: Best F1 score, AUPR (Area Under the Precision-Recall Curve), Affiliation F1  (Huet et al., 2022), PATE  (Ghorbani et al., 2024), and AUROC (Area Under the Receiver Operating Characteristic Curve) for TSAD performance evaluation. This combination of metrics provides a comprehensive, fair and robust TSAD evaluation. Detailed descriptions of evaluation metrics are provided in Appendix E.3.
4.4. Implementation Details
@@
Our experiments are conducted on a server with an NVIDIA A40 GPU, 13 CPU cores, and 250GB RAM. We employ the unsupervised TSAD network architecture and training scheme from CARLA as our default TSAD model to perform GenIAS-enhanced detection. GenIAS is implemented in Python using PyTorch and trained using the Adam optimizer with a learning rate of $10^{-4}$ for 1000 epochs. The default window size $T$ is set to 200, and latent dimension size $L$ is set to 100 and 50, for MTS and UTS respectively. The hyperparameters for the total loss are $\alpha=1.0$, $\beta=0.1$, $\gamma=0.0/0.01$ (for UTS/MTS), and $\zeta=0.1$, which are fixed for all datasets. We adopt the baseline results that are produced using the original implementation and maintain their default hyperparameter settings. Due to the page limit, the detailed implementation details and hyperparameter settings are provided in Appendix F.
@@
Table 2. Anomaly generation quality of GenIAS and US-I baselines in ARP, with the best model in bold and the second-best model underlined, where CAP abbreviates CutAddPaste.
@@ 4.5. Generation Performance
We first examine the anomaly generation quality of GenIAS and the competing methods that employ anomaly injection (i.e., the US-I category, see section 4.2) in terms of ARP and EDI, as discussed in Section 4.3. We report the results of ARP and EDI in Table 2. GenIAS achieves the highest average ARP ranking, suggesting it generates overall the most realistic anomalies across all datasets. The realism of the anomaly distribution is essential for providing anomaly-informed supervision in the absence of labelled anomalies. Similarly for EDI, GenIAS also achieves the highest average ranking, indicating good diversity in its generated anomalies while sufficiently illustrating the actual anomalies. This diversity implies that the generated anomalies can also lie beyond the original anomaly distribution, which can be beneficial for learning a more stringent normal class boundary, further improving TSAD performance, provided that they do not interfere with the identification of the normal region. To achieve this, GenIAS treats creating informative but not excessive diversity as a key consideration through its novel perturbation loss.

### v2

@@ 3.2 GenIAS’s Joint Learning
Compact KL Divergence for Distinguishability. The standard KL divergence loss is mainly used to regularize the latent space by aligning the approximate posterior distribution $q(z|\mathbf{X})$, parameterized by $\mu$ and $\sigma^{2}$, with a standard Gaussian prior $p(z)$, defined as $\mathcal{N}(0,1)$. However, this does not explicitly control the spread of the normal latent distribution, which can become overly diffuse and increase overlap between generated anomalies and normal samples. We address this by introducing a compact KL divergence loss with a tunable prior variance $\sigma_{\text{prior}}^{2}$:
$\mathcal{L}_{\text{comp-KL}}=-\frac{1}{2}\sum_{i=1}^{|\mathbf{\mathcal{D}}|}\sum_{j=1}^{L}\Big[1+\log(\sigma^{2}_{ij})-\mu_{ij}^{2}-\frac{\sigma^{2}_{ij}}{\sigma_{\text{prior}}^{2}}+2\log(\sigma_{\text{prior}})\Big],$ (2)
@@
Where $\mu_{ij}$ and $\sigma^{2}_{ij}$ are the mean and variance of the latent distribution for the $j$-th dimension of the $i$-th sample, and $\sigma_{\text{prior}}<1$ is the prior standard deviation. We show in Lemma 3.1 that introducing the tunable prior variance $\sigma_{\text{prior}}$ leads to tighter latent representations.
Lemma 3.1 (Compact Latent Space in Compact KL Loss).
Consider the KL divergence between the estimated posterior and the prior distributions for a latent dimension $j$. If the compact KL loss (Eq.  2) is minimized (or its gradient dominates locally) with a prior variance $\sigma^{2}_{\text{prior}}<1$, then for each latent variable $z_{j}$ the optimal posterior variance satisfies $\sigma^{2}_{j}=\sigma^{2}_{\text{prior}}$. Consequently, the posterior variance of each latent variable is driven toward the prior variance $\sigma^{2}_{\text{prior}}$, thereby imposing a compact spread in the latent space.
The full proof is provided in D.1. We now show in the following proposition that limiting the latent variance improves the distinguishability between generated anomalous and normal latent distributions under variance scaling perturbations.
Proposition 3.2 (Compactness leads to greater separation in the latent space).
@@
The proof is provided in D.3. The increased KL divergence under lower variance implies statistically improved distinguishability. We also illustrate in Fig. 8 in the Appendix how our compact KL loss leads to the separation of reconstruction error in the raw space.
Perturbation Loss for Informative Deviation. We define a perturbation loss to control the magnitude of deviation introduced by latent perturbations composed of two complementary terms.
The first term is a triplet-style margin loss that enforces a minimum separation between the original sample $\mathbf{X}$ and the generated anomalous sample $\widetilde{\mathbf{X}}$, relative to the reconstructed normal sample $\hat{\mathbf{X}}$. This term ensures that generated anomalies deviate sufficiently from the original time series by pushing anomalous windows away from normal windows in the raw time series space with a minimum margin $\delta_{\text{min}}$.
[…]
$\mathcal{L}_{\text{perturb}}=\;$ $\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\Big[\max\big(d(\mathbf{X},\hat{\mathbf{X}})-d(\mathbf{X},\tilde{\mathbf{X}})+\delta_{\text{min}},0\big)\Big]$
$+\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\max\big(d(\mathbf{X},\tilde{\mathbf{X}})-\delta_{\text{max}},0\big)$ (3)
These terms constrain the perturbation magnitude to avoid overly obvious deviations while ensuring anomalies remain distinguishable.
Zero-Perturbation Loss for Zero-valued Dimensions. Time series windows may contain dimensions that are entirely zero, corresponding to inactive or sparse signals. Directly applying perturbations to such dimensions can result in unrealistic anomalies. We introduce a zero-perturbation loss that regularizes deviations in zero-valued dimensions. Specifically, the loss penalizes large deviations in dimensions where the original signal is zero, encouraging perturbations to remain small and plausible while preserving the structural characteristics of these dimensions. The loss is defined as:
$\mathcal{L}_{\text{zero-perturb}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\big(\frac{1}{T\cdot|\mathcal{Z}|}\sum_{t=1}^{T}\sum_{d\in\mathcal{Z}}\big(d(\mathbf{X},\tilde{\mathbf{X}})+1)^{-1}\big),$ (4)
where $\mathcal{Z}$ denotes the set of zero-valued dimensions in $\mathbf{X}$. This term discourages unrealistic perturbations in zero-valued channels while allowing controlled deviations when appropriate.
Reconstruction Loss for Learning Normal. The reconstruction loss is widely used for learning latent representation and is defined as Mean Squared Error (MSE) between an input sample $\mathbf{X}$ and its reconstruction $\hat{\mathbf{X}}$:
$\mathcal{L}_{\text{recon}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\frac{1}{T\cdot D}\sum_{t=1}^{T}\sum_{d=1}^{D}\big(\mathbf{X}[t,d]-\hat{\mathbf{X}}[t,d]\big)^{2},$ (5)
which is included to capture the temporal properties of normal samples as the basis for perturbation.
Overall Objective. The overall objective for GenIAS joint learning can be defined as:
@@
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{comp-KL}},$ (6)
@@
where $\alpha$, $\beta$, $\gamma$, and $\zeta$ are the coefficients for the reconstruction, perturbation, zero-perturbation, and KL divergence losses, respectively.
@@
Figure 3: Effect of deviation-based patching thresholds on generated anomalies across SMAP and Yahoo datasets.
3.3 Deviation Patching Approach
As discussed in Sec.3.1, directly using reconstructed anomalous windows has practical limitations. We therefore introduce a deviation-based patching strategy to selectively inject anomalous components while preserving normal context. Specifically, patching is performed independently for each dimension $d$ according to $\mathbf{X}_{\text{patched},d}=\widetilde{\mathbf{X}}_{d}$ if $\lVert\mathbf{X}_{d}-\widetilde{\mathbf{X}}_{d}\rVert^{2}>\tau\cdot(\max(\mathbf{X}_{d})-\min(\mathbf{X}_{d}))$, otherwise $\mathbf{X}_{d}$, where $\tau$ is a scaling factor controlling the sensitivity of anomaly injection. This strategy replaces entire variable trajectories only when a non-trivial deviation is detected, preserving normal patterns in unaffected dimensions. We show its effect in Fig. 3. The complete procedure is summarized in Algo 2.
4 Experiments
@@
We extensively evaluate GenIAS using a diverse set of datasets and baselines in two parts: (1) Generation Evaluation, comparing GenIAS with other anomaly injection models to assess its capability to produce realistic and diverse anomalies; and (2) Case Study – TSAD Evaluation, benchmarking GenIAS against seventeen competing methods using nine real-world MTS and UTS datasets. To assess its impact on TSAD, we replace CARLA’s anomaly generator with GenIAS while keeping the rest of the architecture unchanged. CARLA is selected as the base detector for its SOTA performance, enabling fair comparison with the strongest injection-based method. This comprehensive evaluation allows for a detailed analysis of GenIAS’s effectiveness in both anomaly generation and TSAD.
Table 1: An overview of dataset statistics.
@@
Benchmark # datasets # dims Train size Test size AR% MSL (Hundman et al., 2018) 27 55 58,317 73,729 10.72% SMAP (Hundman et al., 2018) 55 25 140,825 444,035 13.13% SMD (Su et al., 2019) 28 38 708,405 708,420 4.16% SWaT (Mathur and Tippenhauer, 2016) 1 51 496,800 449,919 12.33% GECCO (Moritz et al., 2018) 1 9 69,260 69,261 1.1% SWAN (Angryk et al., 2020) 1 38 60,000 60,000 32.6% UCR (Wu and Keogh, 2021) 250 1 2,238,349 6,143,541 0.6% Yahoo-A1 (Laptev et al., 2015) 67 1 46,667 46,717 1.76% KPI (1) 29 1 1,048,576 2,918,847 1.87%
4.1 Benchmark Datasets
@@
GenIAS is thoroughly evaluated on nine widely used real-world TSAD benchmark datasets, with key statistics reported in Table 1. The selection includes six MTS and three UTS datasets, covering diverse anomaly types and ratios to ensure a comprehensive comparison against baselines. Detailed dataset descriptions are provided in Appendix A.
4.2 Baselines
@@
GenIAS is compared with seventeen competitive baseline methods that leverage a wide range of anomaly detection models, which can be categorized into four broad categories: classic anomaly detection methods, positive-unlabeled (P-U) methods, unsupervised deep TSAD methods that do not leverage anomaly generation (US-WI), and unsupervised TSAD methods that employ anomaly generation (US-I).
@@
Specifically, three classic general AD methods— OCSVM (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008)—to demonstrate the advantages of deep learning TSAD approaches. LSTM-VAE (Park et al., 2018) is included to assess the impact of incorporating labeled data in TSAD. US-WI includes a range of unsupervised TSAD baselines that do not use anomaly generation. Methods such as Donut (Xu et al., 2018), OmniAnomaly (Su et al., 2019), AnomalyTransformer (Xu et al., 2021), and TranAD (Tuli et al., 2022a) rely on reconstruction error for anomaly detection. THOC (Shen et al., 2020b) and TimesNet (Wu et al., 2023) approach TSAD as a forecasting task, while MTAD-GAT (Zhao et al., 2020) incorporates graph attention to model temporal and multivariate dependencies. Representation learning methods like TS2Vec (Yue et al., 2022) and DCdetector (Yang et al., 2023) aim to learn robust, anomaly-discriminative embeddings. US-I, the most relevant category for GenIAS, includes unsupervised TSAD methods that use anomaly generation. We compare GenIAS with four recent models in this category: NCAD (Carmona et al., 2022), CutAddPaste (Wang et al., 2024), COUTA (Xu et al., 2024), and CARLA (Darban et al., 2025).
4.3 Evaluation Metrics
@@
We adopt seven metrics to evaluate GenIAS’s performance in terms of generation quality and TSAD performance.
@@
Generation quality. To quantitatively examine anomaly generation quality, we introduce two key metrics: Anomalous Representation Proximity (ARP), which assesses realism, and Entropy-Based Diversity Index (EDI), which characterizes the diversity, motivated by the idea in AnomalyDiffusion (Hu et al., 2024). Due to space constraints, definition and further discussion are in Appendix E. In general, ARP is the primary metric, reflecting how closely generated anomalies resemble actual ones, while EDI is complementary to ARP by assessing diversity only when realism is satisfied.
@@
TSAD evaluation. We consider five widely-used, complementary metrics to evaluate TSAD performance: F1 score, AUPR (Area Under the Precision-Recall Curve), Affiliation F1  (Huet et al., 2022), PATE  (Ghorbani et al., 2024), and AUROC (Area Under the Receiver Operating Characteristic Curve). A combination of metrics offers a comprehensive, fair, and robust assessment. See Appendix E.3 for details.
4.4 Implementation Details
@@
Experiments are run on a server with an NVIDIA A40 GPU, 13 CPU cores, and 250GB RAM. We adopt the unsupervised TSAD architecture and training scheme from CARLA as the default model for GenIAS-enhanced detection. GenIAS is implemented in Python using PyTorch and trained using the Adam optimizer with a learning rate of $10^{-4}$ for 1000 epochs. The default window size $T$ is set to 200, and the latent dimension size $L$ is set to 100 and 50, for MTS and UTS, respectively. The hyperparameters for the total loss are $\alpha=1.0$, $\beta=0.1$, $\gamma=0.0/0.01$ (for UTS/MTS), and $\zeta=0.1$, which are fixed for all datasets. We use baseline results from their original implementations with default hyperparameters. Detailed implementation and hyperparameter settings are provided in Appendix F.
4.5 Generation Performance
@@
Table 2: Evaluation results of anomaly generation quality for GenIAS and US-I baselines on the ARP metric, with the best in bold and second-best underlined (CAP: CutAddPaste).
@@
We first examine the anomaly generation quality of GenIAS and the competing methods that employ anomaly injection (i.e., the US-I category, see section 4.2) in terms of ARP and EDI (Section 4.3). We report the results of ARP and EDI in Table 2. GenIAS achieves the highest average ARP ranking, suggesting it generates, overall, the most realistic anomalies across all datasets. The realism of the anomaly distribution is essential for providing anomaly-informed supervision in the absence of labeled anomalies. Similarly, for EDI, GenIAS also achieves the highest average ranking, indicating good diversity in its generated anomalies while sufficiently illustrating the actual anomalies. This diversity implies that the generated anomalies can also lie beyond the original anomaly distribution, which can be beneficial for learning a stricter normal class boundary, further improving TSAD performance, provided that they do not interfere with the identification of the normal region. To achieve this, GenIAS treats creating informative but not excessive diversity as a key consideration through its novel perturbation loss.
4.6 Visualization
Figure 4: Time series injection methods on the Yahoo dataset. Real anomalies are shown in blue, and synthetic anomalies in orange.